One hundred theater tickets were sold. Orchestra seats cost $75 and balcony seats cost $40. The total receipts from ticket sales were $5,400. How many of each were sold?
Answer by mananth(16946)

x= Number of Orchestra tickets
y= Number of Balcony tickets

1 x + 1 y = 100 .............1
Total value
75 x + 40 y = 5400 .............2
Eliminate y
multiply (1)by -40
Multiply (2) by 1
-40 x -40 y = -4000
75 x + 40 y = 5400
Add the two equations
35 x = 1400
/ 35
x = 40
plug value of x in (1)
1 x + 1 y = 100
40 + y = 100
y = 100 -40
y = 60
y = 60
x= 40 Number of Orchestra tickets
y= 60 Number of Balcony tickets
m.ananth@hotmail.ca